# Open X-Embodiment: Robotic Learning Datasets and RT-X Models

### v5

@@ I Introduction
00footnotetext: Allen Institute for AI; Arizona State University; California Institute of Technology; Carnegie Mellon University; Columbia University; EPFL; ETH Zürich; Flexiv Robotics; Georgia Institute of Technology; German Aerospace Center; Google DeepMind; Google Research; IO-AI TECH; Imperial College London; Intrinsic LLC; Istituto Italiano di Tecnologia; Korea Advanced Institute of Science & Technology; Max Planck Institute; Meta AI; Mila Quebec; NVIDIA; New York University; Princeton University; Queensland University of Technology; RIKEN; Shanghai Jiao Tong University; Stanford University; Technische Universität Darmstadt; The University of Texas at Austin; The University of Tokyo; Toyota Research Institute; University of California, Berkeley; University of California, Davis; University of California, San Diego; University of Edinburgh; University of Freiburg; University of Illinois Urbana-Champaign; University of Michigan; University of Montreal; University of Pennsylvania; University of Southern California; University of Technology, Nuremberg; University of Texas at Austin; University of WashingtonA central lesson from advances in machine learning and artificial intelligence is that large-scale learning from diverse datasets can enable capable AI systems by providing for general-purpose pretrained models. In fact, large-scale general-purpose models typically trained on large and diverse datasets can often outperform their narrowly targeted counterparts trained on smaller but more task-specific data. For instance, open-vocab classifiers (e.g., CLIP [1]) trained on large datasets scraped from the web tend to outperform fixed-vocabulary models trained on more limited datasets, and large language models [2, 3] trained on massive text corpora tend to outperform systems that are only trained on narrow task-specific datasets. Increasingly, the most effective way to tackle a given narrow task (e.g., in vision or NLP) is to adapt a general-purpose model. However, these lessons are difficult to apply in robotics: any single robotic domain might be too narrow, and while computer vision and NLP can leverage large datasets sourced from the web, comparably large and broad datasets for robotic interaction are hard to come by. Even the largest data collection efforts still end up with datasets that are a fraction of the size and diversity of benchmark datasets in vision (5-18M) [4, 5] and NLP (1.5B-4.5B) [6, 7]. More importantly, such datasets are often still narrow along some axes of variation, either focusing on a single environment, a single set of objects, or a narrow range of tasks. How can we overcome these challenges in robotics and move the field of robotic learning toward large data regime that has been so successful in other domains?

### v6

@@ I Introduction
00footnotetext: 1Allen Institute for AI; 2Arizona State University; 3California Institute of Technology; 4Carnegie Mellon University; 5Columbia University; 6EPFL; 7ETH Zürich; 8Flexiv Robotics; 9Georgia Institute of Technology; 10German Aerospace Center; 11Google DeepMind; 12Google Research; 13IO-AI TECH; 14Imperial College London; 15Intrinsic LLC; 16Istituto Italiano di Tecnologia; 17Korea Advanced Institute of Science & Technology; 18Max Planck Institute; 19Meta AI; 20Mila Quebec; 21NVIDIA; 22New York University; 23Princeton University; 24Queensland University of Technology; 25RIKEN; 26Shanghai Jiao Tong University; 27Stanford University; 28Technische Universität Darmstadt; 29The University of Texas at Austin; 30The University of Tokyo; 31Toyota Research Institute; 32University of California, Berkeley; 33University of California, Davis; 34University of California, San Diego; 35University of Edinburgh; 36University of Freiburg; 37University of Illinois Urbana-Champaign; 38University of Michigan; 39University of Montreal; 40University of Pennsylvania; 41University of Southern California; 42University of Technology, Nuremberg; 43University of Texas at Austin; 44University of WashingtonA central lesson from advances in machine learning and artificial intelligence is that large-scale learning from diverse datasets can enable capable AI systems by providing for general-purpose pretrained models. In fact, large-scale general-purpose models typically trained on large and diverse datasets can often outperform their narrowly targeted counterparts trained on smaller but more task-specific data. For instance, open-vocab classifiers (e.g., CLIP [1]) trained on large datasets scraped from the web tend to outperform fixed-vocabulary models trained on more limited datasets, and large language models [2, 3] trained on massive text corpora tend to outperform systems that are only trained on narrow task-specific datasets. Increasingly, the most effective way to tackle a given narrow task (e.g., in vision or NLP) is to adapt a general-purpose model. However, these lessons are difficult to apply in robotics: any single robotic domain might be too narrow, and while computer vision and NLP can leverage large datasets sourced from the web, comparably large and broad datasets for robotic interaction are hard to come by. Even the largest data collection efforts still end up with datasets that are a fraction of the size and diversity of benchmark datasets in vision (5-18M) [4, 5] and NLP (1.5B-4.5B) [6, 7]. More importantly, such datasets are often still narrow along some axes of variation, either focusing on a single environment, a single set of objects, or a narrow range of tasks. How can we overcome these challenges in robotics and move the field of robotic learning toward large data regime that has been so successful in other domains?